Status Report

Hubble Space Telescope Daily Report #3064 – 26 Feb 2002

By SpaceRef Editor
February 26, 2002
Filed under , ,



PERIOD COVERED: 0000Z (UTC) 02/25/02 – 0000Z (UTC) 02/26/02

Daily Status Report as of 057/0000Z


1.1 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the darks. There was no anomalous activity.

1.2 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3)

The WF/PC-2 was used obtain three dark frames every day to provide
data for monitoring and characterizing the evolution of hot pixels. No
problems were encountered.

1.3 Completed Three Sets of STIS/CCD 9088 (Next Generation Spectral
Library of Stars)

The Space Telescope Imaging Spectrograph (CCD) was used to produce
a “Next Generation” Spectral Library of 600 stars for use in modeling the
integrated light of galaxies and clusters by using the low dispersion UV
and optical gratings of STIS. The library will be roughly equally divided
among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling
the entire HR-diagram in each bin. Such a library will surpass all extant
compilations and have lasting archival value, well into the Next Generation
Space Telescope era. No problems occurred.

1.4 Completed Five Sets of WF/PC-2 9318 (POMS Test Proposal: WFII
Parallel Archive Proposal Continuation)

The WF/PC-2 was used to perform the generic target version of the
WFPC2 Archival Pure Parallel program. The program was used to take
parallel images of random areas of the sky, following the recommendations
of the Parallels Working Group. There were no reported problems.

1.5 Completed Two Sets of STIS/CCD 9317 (Pure Parallel Imaging Program:
Cycle 10)

The Space Telescope Imaging Spectrograph (CCD) was used to perform
the default archival pure parallel program for STIS during cycle 10. There
were no reported problems.

1.6 Completed Five Sets of WF/PC-2 8935 (Cycle 10 Standard Darks)

The WF/PC-2 was used to obtain dark frames every week in order to
provide data for the ongoing calibration of the CCD dark current rate, and
to monitor and characterize the evolution of hot pixels. Over an extended
period these data will also provide a monitor of radiation damage to the
CCDs. The proposal completed with no reported problems.

1.7 Completed STIS/MA2 8590 (UV Imaging and Spectroscopy of Luminous
Blue Compact Galaxies from z=0 to z=1)

The Space Telescope Imaging Spectrograph (MA2) was used to observe
two well-defined samples of low-mass starburst galaxies, one in the local
universe {z<0.1} and another at intermediate redshifts {0.2 < z < 0.7}. Both samples show optical sizes, morphologies, emission line widths, and luminosities comparable to those of LBGs at z=3, and are therefore probably the best local analogs and testbeds for further study of LBGs. Our main goals are to: {1} explore the morphologies, surface brightness distributions, and half-light radii of nearby starforming galaxies in the FUV, near Ly-alpha; {2} search for systematic differences among UV, optical, and near-IR morphologies and structural parameters; {3} investigate the intrinsic emission and absorption spectra near Ly-alpha of starbursting dwarf galaxies, with special attention to Ly- alpha profiles and interstellar and stellar photospheric absorption from Si II, O I, C II, Si IV, and C IV; {4} measure their FUV-optical colors and dust extinction properties; and {5} test the hypothesis that low-mass starbursts are the local counterparts of LBGs. The observation completed normally.

1.8 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor)

The Space Telescope Imaging Spectrograph (MA2) was used to perform
daily monitoring of the NUV MAMA detector dark noise in order to monitor
the effects of thermal changes on the NUV dark rate. No problems occurred.

1.9 Completed STIS/CCD/MA1 9054 (Star Formation Triggers and Chemical
Reprocessing in I Zw 18)

The Space Telescope Imaging Spectrograph (CCD and MA1) was used to
observe I Zw 18, a blue compact galaxy {BCD} with the smallest known
abundance of heavy elements yet observed. No anomalies were reported.

1.10 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor)

The WF/PC-2 was used to calibrate the internal monitor, to be run
weekly to monitor the health of the cameras. No problems were encountered.

1.11 Completed STIS/CCD 8903 (Bias Monitor – Part 1)

The Space Telescope Imaging Spectrograph (CCD) was used to monitor
the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at
gain = 4 to build up high-S/N superbiases and track the evolution of hot
columns. There were no problems.

1.12 Completed WF/PC-2 9180 (Gamma-ray Burst Progenitors: Probing Their

The WF/PC-2 was used to perform a target of opportunity observation
of gamma ray burster (GRB), GRB-020124. GRB astronomy is a field maturing
at a phenomenal rate. Three important new observational and theoretical
discoveries, formulated over the last twelve months, allow the proposer to
address new, and in many cases, more sophisticated questions than could
have been posed previously. These developments: the discovery of X-ray
lines in GRB 991216; the observation that N_H as deduced from X-ray
afterglow are one to two orders of magnitude larger than the dust
extinction inferred from optical afterglow; and the growing realization
that the afterglow emission may exhibit features of dust echoes, appear to
offer unexpected and new diagnostics that will directly inform us about the
progenitor, the circum-progenitor material and the immediate interstellar
environs. There were no reported problems.


2.1 Guide Star Acquisitions:

Scheduled Acquisitions: 5

Successful: 5

Scheduled Re-acquisitions: 9

Successful: 9

2.2 FHST Updates:

Scheduled: 12

Successful: 9

As detailed in HSTAR 8509, the tracker 2 and 3 full maneuver
updates at 057/011102Z and at 057/011347Z failed on tracker 3. The
following acquisition at 057/011705Z, however, was successful. Then, the
tracker 2 and 3 roll delay update at 057/023508Z also failed due to tracker 3.

2.3 Operations Notes:

Using ROP DR-18-A, the CCS limit for SESBSLD was modified twice.

The flight operations team transferred to the GSFC STOCC at
056/1530Z in preparation for Servicing Mission 3B.

The ephemeris table was loaded at 057/0023Z per ROP DF-07A.

SSR-1 EDAC errors were cleared at 057/0258Z as directed by ROP SR-1A.


Servicing Mission 3B, currently scheduled for February 28.

Continuation of normal science observations and calibrations.

SpaceRef staff editor.